OpenStudy (anonymous):

Calculus 2 Does the following (sequence) converge? if it does , what is the limit or explain why it converges, if it diverges explain why. 0.2, 0.24, 0.246, 0.2468, 0.24680, 0.246802

OpenStudy (anonymous):

i would say if the pattern continues it converges to \[\overline{.24680}\]

OpenStudy (anonymous):

you can convert this repeating decimal to a fraction in the usual method

OpenStudy (anonymous):

2/100 +24/100..etc?

OpenStudy (anonymous):

no not for this one, you need \[\frac{24680}{100,000}+\frac{24680}{100,000^2}+...\]

OpenStudy (anonymous):

or use the old trick of writing \[x=\overline{.24680}\\ 100000x=24680\overline{.24680}\\ 99999x=24680\\ x=\frac{24680}{99999}\]

OpenStudy (anonymous):

if it converges, what is the limit?

OpenStudy (anonymous):

that is the limit, the fraction i wrote above

OpenStudy (anonymous):

I have never heard of this old trick though

OpenStudy (anonymous):

We just took sequinces and series recently and I am not so good at them :(

OpenStudy (anonymous):

then you can write it as \[24680(\frac{1}{10^5}+\frac{1}{10^{10}}+...\] and sum the geometric sequence same answer

OpenStudy (anonymous):

Alright! thank you very much

OpenStudy (anonymous):

Appreciate it

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

\[\sum_{n=2}^{\infty} e ^{1/n}/n ^{2}\] does this series converge or diverge?